OpenStudy (anonymous):

For f(x)=integral 1/sqrt(4+e^t)dt from 0 to lnx, with x>0, find (f^-1)'(0)

OpenStudy (anonymous):

so is it \[f(x)=\int_0^{\ln x}\dfrac{dt}{4+e^t},x>0\\f^{-1}(0)=?\]

OpenStudy (anonymous):

or so is it \[f{'}(0)=?\]

OpenStudy (anonymous):

it is (f^-1)'(0)= It's asking for the derivative of the inverse of f at 0

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

it's asking for the derivative of f^-1 at x = 0

OpenStudy (anonymous):

yes, that is correct

OpenStudy (anonymous):

do you know the formula?

OpenStudy (anonymous):

I'm not sure what formula to use

OpenStudy (anonymous):

\[\frac{d}{dx} f^{-1}(x) = \frac{1}{f'(f^{-1}(x))}\]

OpenStudy (anonymous):

two things you need to find d/dx f(x) and f^-1(0)

OpenStudy (anonymous):

I believe the derivative is 1/xsqrt(4+x). Now I have to find the inverse at 0?

OpenStudy (anonymous):

f^-1(0) means 0 = f(x)

OpenStudy (anonymous):

do you know how to find 0 = f(x) ?

OpenStudy (anonymous):

would i just put 1/sqrt(4+e^t)=0?

OpenStudy (anonymous):

no, its